OpenStudy (anonymous):

The area of a rectangle is 65m^2, and the length of the rectangle is 3m less than twice the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

@tester97 try this lol

OpenStudy (anonymous):

@Euler271 dude... just in time.

OpenStudy (anonymous):

lol hey

OpenStudy (anonymous):

so area is L * W = length * width the equation says L = 2W - 3 (3 less than twice the width) so L * W = 65 plug in L = 2W - 3 (2W - 3) * W = 65 solve for W

OpenStudy (anonymous):

-5, 13/2

OpenStudy (anonymous):

yes. and since -5 doesn't make sense, we know the answer is W = 13/2 you can plug that W into L = 2W - 3 to find L

OpenStudy (anonymous):

l=2(13/2)-3? l=10? :3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i was about to make food. ill be on later. im working from home today :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

up to you. sometimes latex is useful

OpenStudy (anonymous):

wanna try it out? i'll tell you if you're doing it right

OpenStudy (anonymous):

latex? o.o I got the answer for that one already

OpenStudy (anonymous):

\[\sqrt{this\ is\ latex}\]

OpenStudy (anonymous):

Oh, I got it. The length of a rectangle is 5ft less than three times the width, and the area of the rectangle is 50ft^2. Find the dimensions of the rectangle. L*W L*W=50 L=5w-3 (5w-3)*W=50 W=3/5 +10/w...? what.

OpenStudy (anonymous):

actually 5 less than 3 times is 3w - 5

OpenStudy (anonymous):

afk making food though

OpenStudy (anonymous):

(3w-5)*w=50?

OpenStudy (anonymous):

-10/3, 5

OpenStudy (anonymous):

yes